OpenStudy (ashley1nonly):

The reaction A→B has been experimentally determined to be second order. The initial rate is 0.0100M/s at an initial concentration of A of 0.300M . What is the initial rate at [A]=0.900M ?

OpenStudy (aaronq):

Have you written the rate law?

OpenStudy (ashley1nonly):

no

OpenStudy (aaronq):

that's step number 1.

OpenStudy (ashley1nonly):

rate=k[A]^2

OpenStudy (aaronq):

good stuff, now find \(k\) with the values initially given in the question

OpenStudy (ashley1nonly):

.0100=k[.900]^2 is this the correct setup

OpenStudy (aaronq):

close, the rate of 0.01 M/s is for \([A]=0.300~M\)

OpenStudy (ashley1nonly):

we trippled the .300 to go to .90 so would i tripple the intial rate to get .300

OpenStudy (aaronq):

that would work if the reaction was linear but it's not, it's second order.

OpenStudy (ashley1nonly):

so how would i find the right answer

OpenStudy (aaronq):

first find \(k\)

OpenStudy (ashley1nonly):

first i find k then that would allow me to plus that in for the next equation so .0100=k[.300]^2 k=.1111 or 1/9 then rate=1/9[.900]^2 .09=rate

OpenStudy (aaronq):

perfect

OpenStudy (ashley1nonly):

Yes thank you so much

OpenStudy (aaronq):

no problem !